OpenStudy (anonymous):

Which is a factor of 6x3y + 6xy – 12x2– 12 ? x2 + 1 x + 2 xy – 1 x2 – 2

jhonyy9 (jhonyy9):

x2+1 so the first is right

OpenStudy (anonymous):

6x3y + 6xy – 12x2– 12 =6xy(x^2+1)-12(x^2+1) =6(xy-2)(x^2+1)